# Some remarks on L-equivalence for cubic fourfolds and hyper-Kähler manifolds

### v1

@@ Abstract
We prove that if two very general cubic fourfolds are L-equivalent then they are isomorphic, and we observe that there exist special cubic fourfolds which are L-equivalent but not isomorphic. As a consequence, we provide examples of hyper-Kähler manifolds which are L-equivalent but not birational. We also provide further examples in support of the fact that L-equivalent hyper-Kähler manifolds should be D-equivalent, as conjectured by Meinsma.
@@ 1.1 L-equivalence
It was conjectured by Kuznetsov–Shinder [kuznetsov2018grothendieck, Conjecture 1.6] that simply connected algebraic varieties that are D-equivalent should be also L-equivalent. By a result of Orlov, if two varieties with ample canonical class or anticanonical class are D-equivalent then they are isomorphic. In particular, they are clearly L-equivalent. On the other hand, when the canonical bundle is trivial, this implication is proved to be false in [meinsma2025counterexamples, Theorem 1.3]. Indeed, there exist examples of projective hyper-Käher manifolds that are D-equivalent but not L-equivalent. The opposite implication is not entirely clear in general, and Meinsma conjectures [meinsma2025counterexamples, Conjecture 1.5] that two L-equivalent projective hyper-Kähler manifolds must be D-equivalent.
@@ 1.2 Results
Such examples are cubic fourfolds $Y,Y^{\prime}\in\mathcal{C}_{20}$ which were studied by Fan-Lai [Fan_Lai:New_rational], they are birational to each other. These cubic fourfolds are also Furier-Mukai (FM) partners, meaning that there is a Furier-Mukai equivalence $\mathcal{A}_{Y}\cong\mathcal{A}_{Y^{\prime}}$ between the Kuznetsov components [kuznetsov2009derived], here $\mathcal{A}_{Y}:=\langle\mathcal{O},\mathcal{O}(1),\mathcal{O}(2)\rangle^{\perp}\subset D^{b}(Y)$.
@@
We now turn to the case of hyper-Kähler manifolds. The first examples of hyper-Kähler manifolds that are L-equivalent and D-equivalent but not birational were constructed by Okawa in [okawa2021example]. Here we give further examples of L-equivalent hyper-Kähler manifolds which are not birational, we expect them to be also D-equivalent in view of [meinsma2025counterexamples, Conjecture 1.5]. Recall that the Fano variety of lines $F(Y)$ of a cubic fourfold $Y$ is hyper-Kähler deformation equivalent to the Hilbert square of a K3 surface [beauville1985variete].
Theorem 1.6 (3.5 and 4.5).
Let $Y\in\mathcal{C}_{20}$ be a general cubic fourfold, then there exists a cubic fourfolds $Y^{\prime}$ such that the Fano varieties of lines $F(Y)$ and $F(Y^{\prime})$ are L-equivalent but not birational.
We also provide some additional evidence for the Meinsma’s conjecture, in some particular cases. If $X$ is hyper-Kähler, we consider the transcendental lattice $T(X)\subset H^{2}(X,\mathbb{Z})$, the smallest sublattice such that $H^{2,0}(X)\subset T(X)\otimes\mathbb{C}$, and the Néron-Severi lattice $\operatorname{NS}(X)\subset H^{2}(X,\mathbb{Z})$ which can be seen as the orthogonal to $\operatorname{NS}(X)$ with respect to the Beauville–Bogomolov–Fujiki form, and it consists of algebraic classes.
Theorem 1.7 (4.6, 4.11, 4.13, 4.14, 4.16).
@@
    If $X$ and $X^{\prime}$ are moduli spaces on the Kuznetsov component of very general cubic fourfolds $Y$ and $Y^{\prime}$, then $Y\simeq Y^{\prime}$ and $X$ is birational to $X^{\prime}$.
@@
    If $X=S^{[n]}$ for n1n-1 a prime power and a K3 surface $S$ with $U\subseteq\operatorname{NS}(S)$ and $\operatorname{rk}\operatorname{NS}(S)\not=4$, then in case $\operatorname{disc}T(X)=\operatorname{disc}T(X^{\prime})=m$ is coprime to 2(n1)2(n-1) we have that $X$ and $X^{\prime}$ are D-equivalent.
@@ Acknowledgements
Simone Billi would like to thank the Matthias Schütt and the Riemann Center for Geometry and Physics of Hannover for hosting and supporting him during the writing of this article.
@@ 4.1.1 The Fano variety of lines
The following gives more examples of projective hyper-Kähler manifolds that are L-equivalent, but not birational. The first such examples were constructed in [okawa2021example], here we present some cases that were not included and that do not have an associate K3 surface (see [hassett2000special]). The previous examples are also known to be D-equivalent, we expect our examples to be D-equivalent as well in view of [meinsma2025counterexamples, Conjecture 1.5].
Proposition 4.5.
Let $Y\in\mathcal{C}_{20}$ be a very general cubic fourfold, and let $Y^{\prime}\in\mathcal{C}_{20}$ be as in 3.4. Then the Fano varieties of lines $F(Y)$ and $F(Y^{\prime})$ are L-equivalent, but not birational.
Proof.
Firstly, we know that $F(Y)$ and $F(Y^{\prime})$ are L-equivalent by 4.4. We also recall that there is a Hodge isometry $H^{4}_{prim}(Y,\mathbb{Z})\cong H^{2}_{prim}(F(Y),\mathbb{Z})(-1)$, where $H^{2}_{prim}(F(Y),\mathbb{Z})$ is the orthogonal of the polarization $H$ of $F(Y)$, and $H^{2}=6$, from [beauville1985variete].
Assume $F(Y)$ is birational to $F(Y^{\prime})$. Then by the Torelli Theorem for hyper-Kähler manifold [markman2011survey, Theorem 1.3] there exists a Hodge isometry $H^{2}(F(Y),\mathbb{Z})\cong H^{2}(F(Y^{\prime}),\mathbb{Z})$, which might a priori not preserve the polarization. The algebraic lattice of a general cubic $Y\in\mathcal{C}_{20}$ is given by
$A(Y)=\langle\eta_{Y},V\rangle\cong\begin{bmatrix}3&4\\ 4&12\end{bmatrix}$
where $V\subset Y$ is a Veronese surface, see [Fan_Lai:New_rational] for example. The algebraic primitive lattice is then given by $A_{prim}(Y)=\eta_{Y}^{\perp A(Y)}=\langle-4\eta_{X}+3V\rangle\cong[60]$. Let $H$ be the polarization of $F(Y)$, then $\operatorname{NS}(F(Y))$ is an overlattice of $\langle H\rangle\oplus\langle\alpha\rangle$ with $H^{2}=6$ and $\alpha^{2}=-60$. Since $\operatorname{disc}T(Y)=\operatorname{disc}T(F(Y))=20$, the index of the inclusion $\langle H\rangle\oplus\langle\alpha\rangle\subset\operatorname{NS}(F(Y))$ must be $3$. This shows that
$\operatorname{NS}(F(Y))=\langle H,\frac{H-\alpha}{3}\rangle\cong\begin{bmatrix}6&2\\ 2&-6\end{bmatrix}$
and one can check that $O(\operatorname{NS}(F(Y))\cong\{\pm\mathrm{id}\}$. In particular, up to composing with $\mathrm{id}$, twe find a Hodge isometry $H^{2}(F(Y),\mathbb{Z})\cong H^{2}(F(Y^{\prime}),\mathbb{Z})$ preserving the polarization and this induces a Hodge isometry $H_{prim}^{4}(Y,\mathbb{Z})\cong H_{prim}^{4}(Y^{\prime},\mathbb{Z})$. However, this is not possible since $Y\not\cong Y^{\prime}$ and we conclude that $F(Y)$ and $F(Y^{\prime})$ are not birational. ∎
@@ 4.1.3 Moduli spaces on the Kuznetsov component
If $v=kv_{0}$ for $k\in\mathbb{Z}$ and $v_{0}$ primitive we get that an element $u\in v^{\perp}\iff u\in v_{0}^{\perp}$, so the same result of 4.9 holds for $v=kv_{0}$, $w=kw_{0}$ with $v_{0},w_{0}$ primitive vectors with the same square.
Now we are ready to state and prove the following proposition in the same fashion of 4.6.
@@
Let $Y$ and $Y^{\prime}$ be two cubic fourfolds and assume that $Y$ is very general. Let $v\in A(\mathcal{A}_{Y})$ and $w\in A(\mathcal{A}_{Y^{\prime}})$ be either primitive with same square or of O’Grady type. Let $\sigma_{Y}\in\operatorname{Stab}^{\dagger}(Y)$ and $\sigma_{Y^{\prime}}\in\operatorname{Stab}^{\dagger}(Y^{\prime})$ be a $v$-generic and a $w$-generic stability condition, respectively. Suppose that $M_{v}(\sigma_{Y})$ and $M_{w}(\sigma_{Y^{\prime}})$ are L-equivalent, then $Y\cong Y^{\prime}$ and $M_{v}(\sigma_{Y})$ is birational to $M_{w}(\sigma_{Y^{\prime}})$.
@@
By 4.8 we get that $T(M_{v}(\sigma_{Y}))$ is Hodge anti-isometric to $T(Y)$ and $T(M_{w}(\sigma_{Y}))$ to $T(Y^{\prime})$. Since $Y$ is very general we know by 2.8 that $\operatorname{End}(T(X))=\mathbb{Z}$. Since $Y$ is very general, we have $T(Y)=H_{prim}^{4}(Y,\mathbb{Z})$ and similarly for $Y^{\prime}$. From the fact that $T(Y)$ and $T(Y^{\prime})$ are isometric of signature $(22,2)$, by 2.7 and 2.9 we get Hodge isometries $T(Y)\simeq T(M_{v}(\sigma_{Y}))(-1)\simeq T(M_{w}(\sigma_{Y^{\prime}}))(-1)\simeq T(Y^{\prime})$. From the first and the last Hodge isometries we get a Hodge isometry $H^{4}(Y,\mathbb{Z})\cong H^{4}(Y^{\prime},\mathbb{Z})$. The Torelli theorem for cubic fourfolds implies then then $Y\simeq Y^{\prime}$. Moreover, we have $A(\mathcal{A}_{Y})\simeq A(\mathcal{A}_{Y^{\prime}})\simeq A_{2}$ and by [AT, Proposition 2.8] we get Hodge isometries $A_{2}^{\perp}=T(\mathcal{A}_{Y})\cong H^{4}_{prim}(Y,\mathbb{Z})(-1)\simeq H^{4}_{prim}(Y^{\prime},\mathbb{Z})(-1)\cong T(\mathcal{A}_{Y^{\prime}})$.
As $v,w\in A_{2}$ share the same square and are either primitive or of O’Grady type we deduce from 4.9, with the fact that $O(A_{2})\to O(D({A_{2}}))$ is surjective, that they have isomorphic embeddings in $A_{2}$. So, we can deduce the following chain of Hodge isometries
@@
If $v$ and $w$ are primitive, then $M_{v}(\sigma_{Y})$ and $M_{w}(\sigma_{Y^{\prime}})$ are of type $K3^{[n]}$ with $n=\frac{v^{2}+2}{2}$. By the previous discussion, the Hodge isometry $T(\mathcal{A}_{Y})\cong T(\mathcal{A}_{Y^{\prime}})$ extends to a Hodge isometry of the Mukai lattices $\widetilde{H}(\mathcal{A}_{Y},\mathbb{Z})\cong\widetilde{H}(\mathcal{A}_{Y^{\prime}},\mathbb{Z})$ restricting to (2) with the opportune identifications, and hence [markman2011survey, Corollary 9.9] yields that $M_{v}(\sigma_{Y})$ is birational to $M_{w}(\sigma_{Y^{\prime}})$. ∎
@@
By [LPZ_OG10, Proposition 6.7] the OG10 manifold $\widetilde{M}_{v}(\sigma)$ for $v$ of O’Grady type provides a compactification of the Lagrangian fibration of twisted intermediate Jacobians. Thus, $\widetilde{M}_{v}(\sigma)$ is birational to the twisted LSV manifold $J^{t}(Y)$ [voisin2016hyper]. In particular, when $Y$ is very general, for this choice of $v$, we have that 4.11 and 4.6 are equivalent.
@@
Suppose that in the hypothesis of 4.11 we ask $Y$ to be very general in $\mathcal{C}_{d}$ for $d$ not divisible by 9. Then $Y$ and $Y^{\prime}$ are FM partners.
@@
However, in general we do not get birationality of the moduli spaces in this case as shown in 4.5.
@@ 4.2.1 Desingularizations of moduli spaces of sheaves on a K3
Since $\operatorname{disc}T(X)=1$ there is a primitive vector $\sigma\in\operatorname{NS}(X)$ such that $\sigma^{2}=-6$ of divisibility $3$ in $H^{2}(X,\mathbb{Z})$. By [FGG:OG10_as_moduli, Theorem 3.13] we have that $X$ is birational to a moduli space of sheaves on a K3 surface. Moreover, we have $\operatorname{rk}T(X)=20>13$ and by 4.2 there is a Hodge isometry $T(X)\cong T(X^{\prime})$ in view of 4.3. In particular, $T(X^{\prime})$ is also unimodular and from this it follows that $\operatorname{NS}(X^{\prime})$ has the same discriminant form as $\operatorname{NS}(X)$ (which is the one of $H^{2}(X,\mathbb{Z})$), hence $\operatorname{NS}(X)$ and $\operatorname{NS}(X^{\prime})$ are in the same genus. Since the two lattices are even indefinite and $\operatorname{rk}\operatorname{NS}(X)\geq l(\operatorname{NS}(X))+2$, we can extend the Hodge isometry $T(X)\cong T(X^{\prime})$ to a Hodge isometry $H^{2}(X,\mathbb{Z})\cong H^{2}(X^{\prime},\mathbb{Z})$ by 2.5. Up to composing it with $\mathrm{id}$, the Hodge isometry can be taken to be monodromy by [onorati_monodromy, Theorem 5.4] and the Torelli Theorem for hyper-Kähler manifolds [markman2011survey, Theorem 1.3] implies that $X$ and $X^{\prime}$ are birational.

### v2

@@ Abstract
We prove that if two very general cubic fourfolds are L-equivalent then they are isomorphic, and we observe that there exist special cubic fourfolds which are L-equivalent but not isomorphic. When the cubic fourfolds are very general in certain Hassett divisors, we prove that if they are L-equivalent then they are also Fourier-Mukai partners. We also provide further examples in support of the fact that L-equivalent hyper-Kähler manifolds should be D-equivalent, as conjectured by Meinsma.
@@ 1.1 L-equivalence
It was conjectured by Kuznetsov–Shinder [kuznetsov2018grothendieck, Conjecture 1.6] that simply connected algebraic varieties that are D-equivalent should be also L-equivalent, independently this was also conjectured by Ito-Miura-Okawa-Ueda [ito2020derived, Problem 1.2] without the assumption of simply connectedness. By a result of Orlov, if two varieties with ample canonical class or anticanonical class are D-equivalent then they are isomorphic. In particular, they are clearly L-equivalent. On the other hand, when the canonical bundle is trivial, this implication is proved to be false in [efimov2018some, ito2020derived, meinsma2025counterexamples] including the simply connected case. Indeed, there exist examples of abelian varieties and projective hyper-Kähler manifolds that are D-equivalent but not L-equivalent. A variant of this conjecture [ito2020derived, Problem 7.2] is still an open case and it was proved to be true for abelian varieties in [caucci2023derived]. The opposite implication is not entirely clear in general, and Meinsma conjectures [meinsma2025counterexamples, Conjecture 1.5] that two L-equivalent projective hyper-Kähler manifolds must be D-equivalent.
@@ 1.2 Results
Such examples are cubic fourfolds $Y,Y^{\prime}\in\mathcal{C}_{20}$ which were studied by Fan-Lai [Fan_Lai:New_rational], they are birational to each other. These cubic fourfolds are also Furier-Mukai (FM) partners, meaning that there is a Furier-Mukai equivalence $\mathcal{A}_{Y}\cong\mathcal{A}_{Y^{\prime}}$ between the Kuznetsov components [kuznetsov2009derived], here $\mathcal{A}_{Y}:=\langle\mathcal{O},\mathcal{O}(1),\mathcal{O}(2)\rangle^{\perp}\subset D^{b}(Y)$. [^4] A similar situation is also expected in the case of cubic fourfolds containing pairs of non-syzygetic cubic scrolls, as in [brooke2024birational]
@@
We now turn to the case of hyper-Kähler manifolds and provide some additional evidence for the Meinsma’s conjecture, in some particular cases. We recall that birational hyper-Kähler manifolds are conjectured to be also D-equivalent [bondal1995semiorthogonal, kawamata2002d], and this is known to be true in case of manifolds of K3[n] type [maulik2025d]. If $X$ is hyper-Kähler, we consider the transcendental lattice $T(X)\subset H^{2}(X,\mathbb{Z})$, the smallest sublattice such that $H^{2,0}(X)\subset T(X)\otimes\mathbb{C}$, and the Néron-Severi lattice $\operatorname{NS}(X)\subset H^{2}(X,\mathbb{Z})$ which can be seen as the orthogonal to $\operatorname{NS}(X)$ with respect to the Beauville–Bogomolov–Fujiki form, and it consists of algebraic classes.
Theorem 1.6 (4.5, 4.10, 4.12, 4.13, 4.14).
@@
    If $X$ and $X^{\prime}$ are moduli spaces on the Kuznetsov component of very general cubic fourfolds $Y$ and $Y^{\prime}$ with Mukai vector of the same square, then $Y\simeq Y^{\prime}$ and $X$ is birational to $X^{\prime}$. In case $X$ and $X^{\prime}$ are of K3[n] type, they are also D-equivalent.
@@
    If $X$ is birational to $S^{[n]}$ for a K3 surface $S$ with $U\subseteq\operatorname{NS}(S)$ and $\operatorname{rk}T(S)\not=4$, then in case $\operatorname{disc}T(X)=\operatorname{disc}T(X^{\prime})=m$ is coprime to 2(n1)2(n-1) we have that $X$ and $X^{\prime}$ are D-equivalent.
@@ Acknowledgements
We are thankful to Lisa Marquand for pointing out a mistake in the previous version and for useful comments, also to Dominique Mattei for bringing the work [maulik2025d] to our attention.
@@
Simone Billi would like to thank Matthias Schütt and the Riemann Center for Geometry and Physics of Hannover for hosting and supporting him during the writing of this article.
@@ 4.1.1 The Fano variety of lines
In particular, if $Y$ and $Y^{\prime}$ are cubic fourfolds as in 3.4 then by 3.5 and 4.4 we get that $F(Y)$ and $F(Y^{\prime})$ are L-equivalent. However, it was already known by [brooke2024cubic, Theorem 4.1] that $F(Y)\cong F(Y^{\prime})$.
@@ 4.1.3 Moduli spaces on the Kuznetsov component
If $v=kv_{0}$ for $k\in\mathbb{Z}$ and $v_{0}$ primitive we get that an element $u\in v^{\perp}\iff u\in v_{0}^{\perp}$, so the same result of 4.8 holds for $v=kv_{0}$, $w=kw_{0}$ with $v_{0},w_{0}$ primitive vectors with the same square.
Now we are ready to state and prove the following proposition in the same fashion of 4.5.
@@
Let $Y$ and $Y^{\prime}$ be two cubic fourfolds and assume that $Y$ is very general. Let $v\in A(\mathcal{A}_{Y})$ and $w\in A(\mathcal{A}_{Y^{\prime}})$ be either primitive with same square or of O’Grady type. Let $\sigma_{Y}\in\operatorname{Stab}^{\dagger}(Y)$ and $\sigma_{Y^{\prime}}\in\operatorname{Stab}^{\dagger}(Y^{\prime})$ be a $v$-generic and a $w$-generic stability condition, respectively. Suppose that $M_{v}(\sigma_{Y})$ and $M_{w}(\sigma_{Y^{\prime}})$ are L-equivalent, then $Y\cong Y^{\prime}$ and $M_{v}(\sigma_{Y})$ is birational to $M_{w}(\sigma_{Y^{\prime}})$. In case $M_{v}(\sigma_{Y})$ and $M_{w}(\sigma_{Y^{\prime}})$ are of K3[n] type, then they are also D-equivalent.
@@
By 4.7 we get that $T(M_{v}(\sigma_{Y}))$ is Hodge anti-isometric to $T(Y)$ and $T(M_{w}(\sigma_{Y}))$ to $T(Y^{\prime})$. Since $Y$ is very general we know by 2.8 that $\operatorname{End}(T(X))=\mathbb{Z}$. Since $Y$ is very general, we have $T(Y)=H_{prim}^{4}(Y,\mathbb{Z})$ and similarly for $Y^{\prime}$. From the fact that $T(Y)$ and $T(Y^{\prime})$ are isometric of signature $(22,2)$, by 2.7 and 2.9 we get Hodge isometries $T(Y)\simeq T(M_{v}(\sigma_{Y}))(-1)\simeq T(M_{w}(\sigma_{Y^{\prime}}))(-1)\simeq T(Y^{\prime})$. From the first and the last Hodge isometries we get a Hodge isometry $H^{4}(Y,\mathbb{Z})\cong H^{4}(Y^{\prime},\mathbb{Z})$. The Torelli theorem for cubic fourfolds implies then then $Y\simeq Y^{\prime}$. Moreover, we have $A(\mathcal{A}_{Y})\simeq A(\mathcal{A}_{Y^{\prime}})\simeq A_{2}$ and by [AT, Proposition 2.8] we get Hodge isometries $A_{2}^{\perp}=T(\mathcal{A}_{Y})\cong H^{4}_{prim}(Y,\mathbb{Z})(-1)\simeq H^{4}_{prim}(Y^{\prime},\mathbb{Z})(-1)\cong T(\mathcal{A}_{Y^{\prime}})$.
As $v,w\in A_{2}$ share the same square and are either primitive or of O’Grady type we deduce from 4.8, with the fact that $O(A_{2})\to O(D({A_{2}}))$ is surjective, that they have isomorphic embeddings in $A_{2}$. So, we can deduce the following chain of Hodge isometries
@@
If $v$ and $w$ are primitive, then $M_{v}(\sigma_{Y})$ and $M_{w}(\sigma_{Y^{\prime}})$ are of type $K3^{[n]}$ with $n=\frac{v^{2}+2}{2}$. By the previous discussion, the Hodge isometry $T(\mathcal{A}_{Y})\cong T(\mathcal{A}_{Y^{\prime}})$ extends to a Hodge isometry of the Mukai lattices $\widetilde{H}(\mathcal{A}_{Y},\mathbb{Z})\cong\widetilde{H}(\mathcal{A}_{Y^{\prime}},\mathbb{Z})$ restricting to (2) with the opportune identifications, and hence [markman2011survey, Corollary 9.9] yields that $M_{v}(\sigma_{Y})$ is birational to $M_{w}(\sigma_{Y^{\prime}})$. By [maulik2025d, Theorem 1.2] they are also D-equivalent. ∎
@@
By [LPZ_OG10, Proposition 6.7] the OG10 manifold $\widetilde{M}_{v}(\sigma)$ for $v$ of O’Grady type provides a compactification of the Lagrangian fibration of twisted intermediate Jacobians. Thus, $\widetilde{M}_{v}(\sigma)$ is birational to the twisted LSV manifold $J^{t}(Y)$ [voisin2016hyper]. In particular, when $Y$ is very general, for this choice of $v$, we have that 4.10 and 4.5 are equivalent.
@@
Suppose that in the hypothesis of 4.10 we ask $Y$ to be very general in $\mathcal{C}_{d}$ for $d$ not divisible by 9. Then $Y$ and $Y^{\prime}$ are FM partners.
@@ 4.2.1 Desingularizations of moduli spaces of sheaves on a K3
Since $\operatorname{disc}T(X)=1$, the primitive vector $\sigma\in A_{2}(-1)\subset\operatorname{NS}(X)$ with $\sigma^{2}=-6$ and of divisibility $3$ in $A_{2}(-1)$ has then divisibility $3$ also in $H^{2}(X,\mathbb{Z})$. By [FGG:OG10_as_moduli, Theorem 3.13] we have that $X$ is birational to a moduli space of sheaves on a K3 surface. Moreover, we have $\operatorname{rk}T(X)=20>13$ and by 4.2 there is a Hodge isometry $T(X)\cong T(X^{\prime})$ in view of 4.3. In particular, $T(X^{\prime})$ is also unimodular and from this it follows that $\operatorname{NS}(X^{\prime})$ has the same discriminant form as $\operatorname{NS}(X)$ (which is the one of $H^{2}(X,\mathbb{Z})$), hence $\operatorname{NS}(X)$ and $\operatorname{NS}(X^{\prime})$ are in the same genus. Since the two lattices are even indefinite and $\operatorname{rk}\operatorname{NS}(X)\geq l(\operatorname{NS}(X))+2$, we can extend the Hodge isometry $T(X)\cong T(X^{\prime})$ to a Hodge isometry $H^{2}(X,\mathbb{Z})\cong H^{2}(X^{\prime},\mathbb{Z})$ by 2.5. Up to composing it with $\mathrm{id}$, the Hodge isometry can be taken to be monodromy by [onorati_monodromy, Theorem 5.4] and the Torelli Theorem for hyper-Kähler manifolds [markman2011survey, Theorem 1.3] implies that $X$ and $X^{\prime}$ are birational.